. Suppose you just have just enough dimes and quarters to pay for a loaf of bread and quart of milk, which cost $3.45. You have a total of 15 dimes and quarters.
Let d = the number of dimes you have.
Let q = the number of quarters you have.
Write a system of equations that models the information given above.
Solve the system for the number of dimes and quarters you have. You may want to solve this by graphing.

Answer by rcmcc(152)

Let d = the number of dimes you have.
Let q = the number of quarters you have.
so to solve this question you must make 2 simple linear equations.
Since D is dimes and Q is quarters D+Q=15 coins is total.
If you have a total of $3.45 you can use the face value of the coins as the second equation of .10D+.25Q=3.45.
so your two equations would be.
D+Q=15
.10D+.25Q=3.45
You can choose to solve for eith dimes or quarters first, lets solve for quarters. First we must eliminate dimes, so we will multiply the top equation by .10 and then subtract it from the bottom equation.
.10D+.10Q=1.5
-.10D+.25Q=3.45
__________________
-.15Q=-1.95
remove the negatives by multipling by -1 then divide out .15 to isolate Q
.15Q/.15=1.95/.15
Q=13
to find dimes substitute q into the first equation
D+Q=15
D+13=15
D=2
so you have 13 quarters and 2 dimes